OpenStudy (sierra_grey):

Not sure how to do this, Fill in the blank to make the two fractions equivalent. 3/5 = ?/25

OpenStudy (phi):

if you write it as \[ \frac{3}{5} = \frac{x}{25} \] you can think of the problem as, what do I multiply the bottom number (the 5) by to get 25 then use that same number times 3 to get the "x" or, (this is algebra), multiply both sides by 25 like this \[ 25\cdot \frac{3}{5} = 25\cdot\frac{x}{25} \\ 25\cdot \frac{3}{5} = x \]